# Electrically-gated laser-induced spin dynamics in magneto-electric iron garnet at room temperature

### v1

@@ Abstract
In addition, apart from the dominant quadratic response, engineered asymmetry in iron garnets can produce a linear ME contribution [17, 18], while naturally occurring inhomogeneous magneto-electric effects could be exploited [19, 20]. These effects open new possibilities for enhancing electric control of spin dynamics in magnetic materials with magnetoelectric properties, making them promising systems for the continued study of magnetoelectric phenomena and laser-induced spin dynamics.
@@ Acknowledgments
The authors thank A. Stupakiewicz, L. Nowak, L. Shelukhin, and Th. Rasing for fruitful discussions, and K. Saeedi Ilkhchy and C. Berkhout for technical support. The work is supported by the European Research Council ERC Grant Agreement No. 101054664 (SPARTACUS) and by the European Union’s Horizon 2020 Research and Innovation Program under Marie Skłodowska-Curie Grant Agreement No. 861300 (COMRAD). L.K. gratefully acknowledges funding from the Radboud Excellence Initiative. A.P.P acknowledges the support of Russian Ministry of Education and Science (Project No. 075-15-2022-1131) and Basis Foundation “Junior Leader” program.
@@
Timur T. Gareev: Investigation (experiment, equal; modeling, supporting); Data Curation; Visualization (equal); Writing – original draft (equal); Writing –- review and editing (equal). Nikolai E. Khokhlov: Investigation (experiment, equal); Methodology (experiment); Visualization (equal); Writing – original draft (equal); Writing –- review and editing (equal). Lukas Körber: Investigation (theory); Software, Writing –- review and editing (equal). Alexander A. Pyatakov: Methodology (theoretical model); Writing –- original draft (equal). Alexey V. Kimel: Conceptualization; Interpretation; Supervision; Writing –- review and editing (equal); Project Administration.

### v2

@@ Abstract
In addition, apart from the dominant quadratic response, engineered asymmetry in iron garnets can produce a linear ME contribution [17, 18], while naturally occurring inhomogeneous magneto-electric effects could be exploited [19, 20]. Moreover, artificial multiferroic heterostructures offer a powerful platform to optimize electric manipulation of the spin dynamics [21], where interfacial strain, spin-exchange, and charge coupling, as well as voltage-controlled magnetic anisotropy [22], or voltage-driven inversion of magnetic anisotropy via strain-mediated coupling [23], enable reversible electric-field control of magnetism. All of this opens new possibilities for enhancing electric control of spin dynamics in magnetic materials with magnetoelectric properties, making them promising systems for the continued study of magnetoelectric phenomena and laser-induced spin dynamics.
@@ Acknowledgments
The authors thank A. Stupakiewicz, L. Nowak, L. Shelukhin, and Th. Rasing for fruitful discussions, and K. Saeedi Ilkhchy and C. Berkhout for technical support. T.T.G., N.E.Kh., A.V.K. acknowledge support by the European Research Council ERC Grant Agreement No. 101054664 (SPARTACUS) and T.T.G., A.V.K. acknowledge support by the European Union’s Horizon 2020 Research and Innovation Program under Marie Skłodowska-Curie Grant Agreement No. 861300 (COMRAD). A.P.P. acknowledges the support of the Russian Science Foundation (grant No. 25-79-30019) for supporting the theoretical study of the magneto-electric effect. L.K. gratefully acknowledges funding from the Radboud Excellence Initiative.
@@
Timur T. Gareev: Investigation (experiment, equal; modeling, supporting); Data Curation; Visualization (equal); Writing – original draft (equal); Writing –- review and editing (equal). Nikolai E. Khokhlov: Investigation (experiment, equal); Methodology (experiment); Visualization (equal); Writing – original draft (equal); Writing –- review and editing (equal). Lukas Körber: Investigation (theory); Software, Writing –- review and editing (equal). Alexander P. Pyatakov: Methodology (theoretical model); Writing –- original draft (equal). Alexey V. Kimel: Conceptualization; Interpretation; Supervision; Writing –- review and editing (equal); Project Administration.